# Single-Source Regular Path Querying in Terms of Linear Algebra

### v1

@@ Abstract.
A given edge-labelled graph two-way regular path queries (2-RPQs) allow one to use regular languages over labelled edges and inverted edges to constraint paths of interest. 2-RPQs are (partially) adopted in different real-world graph analysis systems and are a part of the GQL ISO standard. But the performance of 2-RPQs on real-world graphs is still a bottleneck for wider adoption. A new single-source 2-RPQ algorithm based on linear algebra is proposed. Utilization of high-performance sparse linear algebra libraries for the algorithm implementation allows one to achieve significant speedup over competitors on real-world data and queries. Our implementation demonstrates better performance on average on Wikidata and the respective query log in comparison with MillenniumDB, FalkorDB, and the algorithm of Diego Arroyuelo et al.
Graph Querying, Regular Path Querying, RPQ, Graph Analysis Algorithms, Sparse Linear Algebra, GraphBLAS
copyright: acmlicensedjournalyear: 2024conference: ; ; booktitle: ccs: Information systems Graph-based database modelsccs: Information systems Query languages for non-relational enginesccs: Theory of computation Regular languagesccs: Mathematics of computing Paths and connectivity problems
@@ 1. Introduction
Language-constrained path querying (Barrett et al., 2000) is a way to search for paths in edge-labeled graphs where constraints are formulated in terms of a formal language. The language restricts the set of accepted paths: the sentence formed by the labels of a path should be in the language. Regular languages are the most popular class of constraints used as navigational queries in graph databases.
@@
Queries that use regular languages to specify constraints are called regular path queries or RPQs. Being proposed in 1989 by Alberto O. Mendelzon and Peter T. Wood (Mendelzon and Wood, 1989), RPQs studying intensively for now. Different extensions of RPQs are studying too: two-way RPQ or 2-RPQ that extend the alphabet of RPQs by the inverse of relationship symbols (Calvanese et al., 2000a), conjunctive RPQ or CRPQ that allows one to check several (parallel) paths (Calvanese et al., 2000b). Regular path constraints and related extensions adopted (partially, in some cases) in a wide range of graph analysis systems and respective query languages, such as Cypher (Francis et al., 2018) and PGQL (van Rest et al., 2016). Moreover, RPQ is a part of ISO standard for a graph query language GQL [^1] ISO/IEC 39075:2024, GQL, https://www.iso.org/standard/76120.html, and core of SPARQL 1.1 RDF query language (201, 2013).
@@
Nevertheless, having a long history of theoretical and applied research, adoption in real-world solutions, RPQs (and related extensions) are still in the focus of research. One of the important direction is an implementation and optimization of RPQ evaluation algorithms (Bonifati et al., 2018) to achieve better performance in real-world cases, and performance-targeted solutions still actively developed (Arroyuelo et al., 2022, 2023b; Ma et al., 2024). Thus, the development of new performant algorithms for RPQs evaluation is an actual challenge that also pointed out by Angela Bonifati et al in (Bonifati et al., 2018). Set of different solutions was proposed to improve RPQ evaluation performance: from special indexes creation (Liu et al., 2021; Arroyuelo et al., 2023b) to utilization of different parallel and distributed models (Nolé and Sartiani, 2016; Faltín et al., 2023; Guo et al., 2021a; Wang et al., 2019).
@@
Sparse linear algebra is a promising way to high-performance graph analysis. This idea is actively developed by GraphBLAS (Kepner et al., 2016) community [^2] GraphBLAS community web page: https://graphblas.org/. Huge amount of graph analysis algorithms, such as PageRank or triangle centralities, can be expressed in terms of linear algebra [^3] Almost full list of graph analysis algorithms in terms of linear algebra: https://github.com/GraphBLAS/GraphBLAS-Pointers. and respective implementations demonstrate promising performance in real-world cases (Pelletier et al., 2021). Even more recently, it was shown that utilization of sparse linear algebra is a way to a high-performance algorithm for more expressive class of queries, Context-Free Path Queries, or CFPQ (Terekhov et al., 2021). But, to our knowledge, there are only a few results on linear algebra based algorithms for RPQ. On the one hand, modern graph database FalkorBD444 Sources of FalkorBD on GitHub: https://github.com/FalkorDB/falkordb (formerly RedisGraph) (Cailliau et al., 2019) is based on SuiteSparse:GraphBLAS (Davis, 2023), reference implementation of the GraphBLAS API, and utilises linear algebra to analyse graphs. This database supports a subset of the Cypher graph query language, including parts of regular constraints. But there is no detailed analysis of respective algorithm. At the same time, linear algebra based RPQ evaluation algorithm was recently proposed by Diego Arroyuelo et al (Arroyuelo et al., 2023a). The proposed solution directly translates regular expressions to expressions over matrices and vectors. Being implemented using sparse matrices and parallel computations, this solution still has performance problems that are evident from the evaluation (Arroyuelo et al., 2023a).
@@
There are BFS-based strategies for RPQ evaluation (e.g. (Koschmieder and Leser, 2012)) [^5] For overview of the primary techniques to evaluate regular path queries, such as relational algebra based and automata based approaches, and ways to optimise it, we refer to the book “Querying Graphs” by Angela Bonifati et al (Bonifati et al., 2018).. At the same time, BFS and its variations can be expressed in terms of linear algebra (Davis, 2019a). It opens a way to develop a linear algebra based algorithm for RPQ evaluation having BFS-like traversal as the core, and we investigate this way.
To summarise, in this work we make the following contributions.
@@
    A new BFS-inspired algorithm (LA 2-RPQ) for single-source (or single-destination as a symmetric case) 2-RPQ is proposed. The algorithm is based on linear algebra: it is expressed in terms of operations over sparse boolean matrices. This fact allows one to utilise high-performance parallel libraries, such as SuiteSparse:GraphBLAS (Davis, 2023), for 2-RPQ evaluation. The correctness of the proposed algorithm is proven.
@@
    Our implementation of the proposed algorithm, based on SuiteSparse:GraphBLAS, is evaluated and compared with other linear algebra based solutions such as FalkorDB and the algorithm proposed by Diego Arroyuelo et al (Arroyuelo et al., 2023a), and with the state-of-the-art graph database MillenniumDB. As a result, we show that our solution is comparable with competitors on real-world dataset Wikidata and the respective query log proposed in MillenniumDB path query challenge (Farias et al., 2023). While for some queries the proposed algorithm is slower than the competitors, it is faster on average: ×$\times 3.9$ for MillenniumDB, ×$\times 19.6$ for FalkorDB, ×$\times 4.2$ for the algorithm of Diego Arroyuelo et al. Moreover, it fits within the 1 minute time limit for all queries, while the other competitors require more time for some queries.
@@ 2. Preliminaries
In this section, we provide the theoretical basics of graph theory and formal language theory required to define the RPQ problem and to describe our solution.
2.1. Basics of graph theory
First we define the graph that we use as a data model.
@@ 2.1. Basics of graph theory
  • $\lambda_{\mathcal{N}}:\Delta\mapsto 2^{\Sigma^{\leftrightarrow}}$ maps transitions to labels;
@@
The set of languages accepted by 2-NFAs coincides with the set of all RLs over $\Sigma^{\leftrightarrow}$. Let $[\![\mathcal{N}]\!]=\mathcal{R}$ where $\mathcal{R}$ is the RL accepted by the 2-NFA $\mathcal{N}$.
In practice special kinds of strings called regular expressions provide a useful facility to define regular languages. The algorithms supplying the transformations between regular expressions, RLs and 2-NFAs are out of the scope of this paper.
Notice that 2-NFA can be seen as a graph $\mathcal{G}_{\mathcal{N}}=\langle Q,\Delta,\Sigma^{\leftrightarrow},\lambda_{\mathcal{N}}\rangle$ equipped with the set of starting states $Q_{S}\subseteq Q$ and the set of final states $Q_{F}\subseteq Q$. Analogously introduce sets $\Delta^{-}$, $\lambda_{\mathcal{N}}^{-}$ and the map $\omega_{\mathcal{N}}=\{a_{1}\cdot a_{2}\cdot...\cdot a_{n}\mid a_{i}\in\lambda_{\mathcal{N}}(\delta_{i})\}$ for the paths $\pi=(\delta_{1},\delta_{2},...,\delta_{n})$ in the graph $\mathcal{G}_{\mathcal{N}}$.
Vice versa a graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ equipped with some $V_{S}\subseteq V$ can bee seen as 2-NFA $\mathcal{N}_{\mathcal{G}}=\langle V,L,E,\lambda_{\mathcal{G}}^{\leftrightarrow},V_{S},V\rangle$. Such 2-NFA accepts the language:
@@
Thus, we can treat evaluation of 2-RPQs with fixed set of starting vertices as an intersection of 2-NFAs.
@@
We have all the required preliminaries to formally state the problem solved by the 2-RPQ algorithm.
@@
Recall two-way regular path query (2-RPQ) a triple $\langle\mathcal{G},\mathcal{R},v_{s}\rangle$ where $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ is a graph, $\mathcal{R}$ is a regular language over the alphabet $L^{\leftrightarrow}$ and $v_{s}\in V$ is a starting node.
Frequent practical cases are queries with fixed starting or final vertex. So, for 2-RPQ $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$ we are interested in an efficient way of evaluating these maps:
  • Single-source reachability:
@@
  • Single-destination reachability:
    $[\![\mathcal{Q}]\!]_{SDR}=\left\{\;u\in V\;\middle|\;\begin{gathered}\exists\text{ 2-path }\pi\text{ in }\mathcal{G}\text{ from }u\text{ to }v_{s}\\ \text{and }\omega^{\leftrightarrow}_{\mathcal{G}}(\pi)\cap\mathcal{R}\neq\operatorname{\varnothing}\\ \end{gathered}\;\right\}$
@@ 3. Linear Algebra, Graphs and Relations
In order to operate with graphs in terms of linear algebra we need to see how algebraic object are connected with set relations and how the relations are connected to graph theory.
@@
  • Sum: $A_{n\times m}\operatorname{\operatorname{\oplus}}B_{n\times m}=(a_{ij}\vee b_{ij})_{n\times m}$.
  • Product: $A_{n\times m}\operatorname{\operatorname{\otimes}}B_{m\times k}=\displaystyle\left(\;\bigvee_{k=1}^{n}a_{ik}\wedge b_{kj}\;\right)_{n\times k}$.
  • Transposition: $A_{n\times m}^{T}=(a_{ji})_{m\times n}$.
  • Zero-matrix: $\operatorname{\operatorname{\mathbf{0}}}_{n\times k}=(0)_{n\times k}$.
  • Mask: $A_{n\times m}\langle C_{n\times m}\rangle=(a_{ij}\wedge c_{ij})_{n\times m}$
  • Complement: $\neg A=(\neg a_{ij})$
@@
Let $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ be an arbitrary graph and $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ an arbitrary 2-NFA. Introduce the sets:
@@
Also note that for the given graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$, $E^{a^{-}}=(E^{a})^{-}$ and $G^{a^{-}}=(G^{a})^{T}$.
4. Single-Source RPQ in Terms of Linear Algebra
@@
output : Matrix $P_{|Q|\times|V|}$
@@ 3 $\{G^{a}\}\leftarrow$ Boolean decomposition of the $\mathcal{G}$ adjacency matrix;
4 $M_{|Q|\times|V|},P_{|Q|\times|V|}\leftarrow\operatorname{\operatorname{\mathbf{0}}}_{|Q|\times|V|}$ ;
5 for $q\in Q_{S}$ do
       $M_{q{v_{s}}}\leftarrow 1$;
        // Set the starting relation
6
7 end for
8while $M\neq\operatorname{\operatorname{\mathbf{0}}}$ do
@@
       $P\leftarrow P\operatorname{\operatorname{\oplus}}M$ ;
        // Accumulate resulting matrix
9

### v2

@@ Abstract.
Two-way regular path queries (2-RPQs) allow one to use regular languages over edges and inverted edges in edge-labelled graph to constrain paths of interest. 2-RPQs are (partially) adopted in different real-world graph analysis systems and have become a part of the GQL ISO standard. However the performance of 2-RPQs on real-world graphs remains a bottleneck for wider adoption. Utilisation of high-performance sparse linear algebra libraries for the algorithm implementation allows one to achieve significant speedup over competitors on real-world data and queries.
We propose a new breadth-first-search-based algorithm that leverages linear algebra for evaluating single-source regular path queries. We integrate it into the LAGraph graph processing algorithm infrastructure and provide in-depth performance comparison on the large real-world knowledge bases. Additionally, we present extensive analysis of its performance across different query types using synthetic data, comparing it with various databases and other linear algebra-based approaches.
@@ 1. Introduction
Language-constrained path querying (Barrett et al., 2000) is a way to search for paths in edge-labeled graphs where constraints are expressed in terms of a formal language. The language restricts the set of valid paths: the sequence of labels along a path must form a sentence belonging to the language. Regular languages are the most popular class of constraints used as navigational queries in graph databases.
@@
Queries that employ regular languages to specify constraints are called regular path queries or RPQs. First introduced in 1989 by Alberto O. Mendelzon and Peter T. Wood (Mendelzon and Wood, 1989), RPQs have been intensively studied. Different extensions of RPQs are studying too: two-way RPQ or 2-RPQ that extends the alphabet of RPQs by the inverse of relationship symbols (Calvanese et al., 2000a), conjunctive RPQ or CRPQ that allows one to check several (parallel) paths (Calvanese et al., 2000b). Regular path constraints and their extensions have been (partially, in some cases) adopted in numerous graph analysis systems and query languages, including Cypher (Francis et al., 2018) and PGQL (van Rest et al., 2016). Moreover, RPQ is part of the ISO standard for the GQL graph query language (ISO/IEC 39075:2024, 2024), and the core of SPARQL 1.1 RDF query language (201, 2013).
@@
Despite their long history of theoretical and applied research, as well as real-world adoption, RPQs (and related extensions) remain in the focus of research. One of the important directions is the implementation and optimization of RPQ evaluation algorithms (Bonifati et al., 2018) to achieve better performance in real world cases, and performance-targeted solutions are still actively developed (Arroyuelo et al., 2022, 2023b; Ma et al., 2024). Thus, designing new, efficient algorithms for RPQ evaluation remains an active challenge, as highlighted by Angela Bonifati et al in (Bonifati et al., 2018). Various approaches have been proposed to enhance RPQ evaluation performance, ranging from specialized indexing techniques (Liu et al., 2021; Arroyuelo et al., 2023b) to parallel and distributed computing models (Nolé and Sartiani, 2016; Faltín et al., 2023; Guo et al., 2021a; Wang et al., 2019).
@@
Sparse linear algebra has emerged as a powerful paradigm for high-performance graph analysis, championed by the GraphBLAS community  (Kepner et al., 2016) [^1] GraphBLAS community web page: https://graphblas.org/. A vast number of graph analysis algorithms, such as PageRank or triangle centralities, can be expressed in terms of linear algebra [^2] An almost complete list of graph analysis algorithms in terms of linear algebra: https://github.com/GraphBLAS/GraphBLAS-Pointers. and the respective implementations demonstrate promising performance in real-world cases (Pelletier et al., 2021). Even more, it has been shown that sparse linear algebra enables a high-performance algorithm for more expressive query classes, such as Context-Free Path Queries (CFPQ) (Terekhov et al., 2021). However, to our knowledge, there are only a few studies on linear-algebra-based RPQ algorithms.
On the one hand, the modern graph database FalkorDB333 Sources of FalkorDB on GitHub: https://github.com/FalkorDB/falkordb (formerly RedisGraph) (Cailliau et al., 2019) is based on SuiteSparse:GraphBLAS (Davis, 2023), reference implementation of the GraphBLAS API, and uses linear algebra for graph analysis. While FalkorDB supports a subset of the Cypher query language, including some regular constraints, there is no detailed analysis of the respective algorithm.
At the same time, the linear-algebra-based RPQ evaluation algorithm was recently proposed by Diego Arroyuelo et al (Arroyuelo et al., 2023a). Despite using sparse matrices and parallel computations, this solution exhibits performance limitations, that are evident from the evaluation (Arroyuelo et al., 2023a).
@@
On the other hand, there are BFS-based strategies for the RPQ evaluation (e.g. (Koschmieder and Leser, 2012)) [^4] For an overview of primary RPQ evaluation techniques, including relational algebra and automata-based approaches, and their optimization, we refer to the book “Querying Graphs” by Angela Bonifati et al (Bonifati et al., 2018).. Notably, BFS and its variants can be expressed using linear algebra (Davis, 2019a), suggesting an opportunity to develop a linear-algebra-based RPQ algorithm with BFS-like traversal at its core. We explore this direction in our work.
To summarize, in this work we make the following contributions.
@@
    A novel BFS-based algorithm (LARPQ) for single-source (or symmetrically, single-destination) 2-RPQ is proposed. The algorithm is based on linear algebra: it is expressed in terms of operations over sparse boolean matrices. This fact allows one to utilize high-performance parallel libraries, such as SuiteSparse:GraphBLAS (Davis, 2023), for 2-RPQ evaluation. The correctness of the proposed algorithm is proven.
@@
    Our implementation of the proposed algorithm, based on SuiteSparse:GraphBLAS, is evaluated and compared with other linear algebra-based solutions such as FalkorDB and the algorithm proposed by Diego Arroyuelo et al (Arroyuelo et al., 2023a), with graph databases Blazegraph and state-of-the-art MillenniumDB. Experimental results on real-world datasets (Wikidata with query logs from the MillenniumDB path query challenge (Farias et al., 2023) and Yago-2S) demonstrate that our solution achieves competitive performance. While occasionally slower on individual queries, our algorithm shows consistent average speedups: 6.8×6.8$\times$6.8 × for the algorithm of Diego Arroyuelo et al, 11.3×11.3$\times$11.3 × for MillenniumDB, 18.9×18.9$\times$18.9 × for FalkorDB, and 16.8×16.8$\times$16.8 × for Blazegraph.
@@ 2. Preliminaries
In this section, we provide the theoretical basics of graph theory and formal language theory required to define the RPQ problem and to describe our solution and the algorithm proposed by Diego Arroyuelo et al.
First we define the edge-labeled graph that we use as a data model.
@@
  • $\lambda_{\mathcal{N}}:\Delta\mapsto 2^{\Sigma^{\leftrightarrow}}$ assigns a set of labels (including inverses) to each transition;
@@
The set of languages accepted by 2-NFAs coincides with the set of all regular languages over $\Sigma^{\leftrightarrow}$. Let $[\![\mathcal{N}]\!]=\mathcal{R}$ where $\mathcal{R}$ is the RL accepted by the 2-NFA $\mathcal{N}$.
Notice that 2-NFA can be seen as a graph $\mathcal{G}_{\mathcal{N}}=\langle Q,\Delta,\Sigma^{\leftrightarrow},\lambda_{\mathcal{N}}\rangle$ equipped with the set of starting states $Q_{S}\subseteq Q$ and the set of final states $Q_{F}\subseteq Q$. Analogously introduce sets $\Delta^{-}$, $\lambda_{\mathcal{N}}^{-}$ and the map $\omega_{\mathcal{N}}(\pi)=\{a_{1}\cdot a_{2}\cdot...\cdot a_{n}\mid a_{i}\in\lambda_{\mathcal{N}}(\delta_{i})\}$ for the paths $\pi=(\delta_{1},\delta_{2},...,\delta_{n})$ in the graph $\mathcal{G}_{\mathcal{N}}$.
Conversely, a graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ equipped with some $V_{S}\subseteq V$ can be seen as 2-NFA $\mathcal{N}_{\mathcal{G}}=\langle V,L,E,\lambda_{\mathcal{G}}^{\leftrightarrow},V_{S},V\rangle$. Such 2-NFA accepts the language:
@@
Thus, we can treat the evaluation of 2-RPQs with a fixed set of starting vertices as an intersection of 2-NFAs.
@@
We have all the necessary preliminaries to formally state the problem solved by the 2-RPQ algorithm.
@@
Recall two-way regular path query (2-RPQ) a 4-tuple $\langle\mathcal{G},\mathcal{R},V_{s},V_{f}\rangle$ where $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ is a graph, $\mathcal{R}$ is a regular language over the alphabet $L^{\leftrightarrow}$, $V_{s}\subseteq V$ is a set of starting nodes, and $V_{f}\subseteq V$ is a set of final nodes.
Frequent practical cases are queries with fixed starting or final vertex. So, for 2-RPQ $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},V_{s},V_{f}\rangle$ we are interested in an efficient way of evaluating these maps:
  • Single-source reachability ($V_{s}=\{v_{s}\},V_{f}=V$ ):
@@
  • Single-destination reachability ($V_{s}=V,V_{f}=\{v_{f}\}$):
    $[\![\mathcal{Q}]\!]_{SDR}=\left\{\;u\in V\;\middle|\;\begin{gathered}\exists\text{ 2-path }\pi\text{ in }\mathcal{G}\text{ from }u\text{ to }v_{f}\\ \text{and }\omega^{\leftrightarrow}_{\mathcal{G}}(\pi)\cap\mathcal{R}\neq\operatorname{\varnothing}\\ \end{gathered}\;\right\}$
For such partial cases we introduce the following short versions: $Q_{SSR}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$ and $Q_{SDR}=\langle\mathcal{G},\mathcal{R},v_{f}\rangle$ respectively.
@@ 3. Linear Algebra, Graphs and Relations
In order to operate with graphs in terms of linear algebra we need to see how algebraic objects are connected with set relations and how the relations are connected to graph theory.
@@
Definition 3.0.
Zero matrix is a rectangular matrix $\operatorname{\operatorname{\mathbf{0}}}_{n\times k}$, such that $\operatorname{\operatorname{\mathbf{0}}}_{ij}=0$ for all valid $i$ and $j$.
Definition 3.0.
For the Boolean matrix $A_{n\times m}$, the matrix $B_{n\times m}$, such that $B_{ij}=\neg A_{ij}$ is a complement matrix for the matrix $A$. We denote complementation of $A$ as $\neg A$.
Definition 3.0.
For the Boolean matrix $A_{n\times m}$, the matrix $B_{m\times n}$, such that $B_{ij}=A_{ji}$ is a transposed matrix for the matrix $A$. We denote transposition of $A$ as $A^{T}$.
Definition 3.0.
For the given Boolean matrices $A_{n\times m}$ and $B_{n\times m}$, the sum $A\oplus B$ is a matrix $C_{n\times m}$ such that matrix $C_{ij}=A_{ij}\vee B_{ij}$.
Definition 3.0.
For the given Boolean matrices $A_{n\times m}$ and $B_{m\times k}$, the product $A\otimes B$ is a matrix $C_{n\times k}$ such that matrix
$C_{ij}=\bigvee_{l=1}^{m}A_{il}\wedge B_{lj}.$
Definition 3.0.
For the given Boolean matrices $A_{n\times m}$ and $B_{n\times m}$, the masking $A\langle B\rangle$ is a matrix $C_{n\times m}$ such that matrix $C_{ij}=A_{ij}\wedge B_{ij}.$
@@
For a given arbitrary graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ and 2-NFA $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ introduce the sets:
@@
Also note that for the given graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$, $E^{a^{-}}=(E^{a})^{T}$ and $G^{a^{-}}=(G^{a})^{T}$.
4. BFS-Based Single-Source RPQ in Terms of Linear Algebra
@@
output : Vector $P^{F}$ of size $1\times|V|$
@@ 3 $\{G^{a}\}\leftarrow$ Boolean decomposition of the $\mathcal{G}$ adjacency matrix;
4 $P_{|Q|\times|V|}\leftarrow\operatorname{\operatorname{\mathbf{0}}}_{|Q|\times|V|}$ ;
5 $M_{|Q|\times|V|}\leftarrow M$ s.t. $M_{qv}=1$ if $q\in Q_{S}$, $v=v_{s}$, otherwise 00;
6 $F_{1\times|Q|}\leftarrow F$ s.t. $F_{1}q=1$ if $q\in Q_{F}$, otherwise 00;
7 while $M\neq\operatorname{\operatorname{\mathbf{0}}}$ do
@@
8       $P\leftarrow P\operatorname{\operatorname{\oplus}}M$